OpenStudy (anonymous):

Help ...The height of a square pyramid is 5 meters shorter than the side of its base. Its volume is 108 cubic meters, how tall is the pyramid V= (1/3)BH B= b^2 H=b-5 108= (1/3)(b-5)(b^2) b^3-5b^2 -324=0

OpenStudy (mayankdevnani):

let the height of square pyramid be `h`, So,base of square pyramid will be `5+h`

OpenStudy (mayankdevnani):

\[\large \bf VOLUME=a^2\times \frac{h}{3}\] \[\large \bf a=base~and~h=height\]

OpenStudy (mayankdevnani):

plug the values,we get \[\large \bf 108=(5+h)^2 \times \frac{h}{3}\]

OpenStudy (mayankdevnani):

and solve for `h`?

OpenStudy (anonymous):

@mayankdevnani i will get a cubic right ? wont that give me more than one value for h?

OpenStudy (mayankdevnani):

yeah !! i think two of them would be -ve or imaginary !

OpenStudy (mayankdevnani):

solve it and then tell me

OpenStudy (anonymous):

okay thank you @mayankdevnani

OpenStudy (mayankdevnani):

your welcome :)